SHORT CASE STUDY     &nb

                                   

SHORT CASE STUDY                                                         

Oriental Motor Assembly Pte Ltd assembles three types of motorcycle at the same factory in Singapore: the 100cc Discover; the 250cc Pulsar and the 1000cc Punch. It sells the motorcycles throughout the world. In response to market pressures Oriental Motors has invested heavily in new manufacturing technology in recent years and, as a result, has significantly reduced the size of its workforce.

Historically, the company has allocated all overhead costs using total direct labour hours, but is now considering introducing activity-based costing (ABC). Oriental Motors accountant has produced the following analysis.

Annual Output

(Units)

Annual Direct labour hours

Selling price ($ per unit)

Raw material cost ($ per unit)

Discover

2000

200,000

4,000

400

Pulsar

1600

220,000

6,000

600

Punch

400

80,000

8,000

900

The three cost drivers that generate overheads are:

  • Deliveries to retailers: the number of deliveries of motorcycles to retail showrooms
  • Set-ups: the number of times the assembly line process is re-set to accommodate a production run of a different type of motorcycle.
  • Purchase orders: the number of purchase orders.

The annual cost driver volumes relating to each activity and for each type of motorcycle are as follows:

Number of deliveries to retailers

Number of set-ups

Number of purchase orders

Discover

100

35

400

Pulsar

80

40

300

Punch

70

25

100

The annual overhead costs relating to these activities are as follows:

Particulars

Amount ($)

Deliveries to retailers

2,400,000

Set-up costs

6,000,000

Purchase orders

3,600,000

All direct labour is paid at $5 per hour. The company holds no stocks. At a board meeting there was some concern over the introduction of activity-based costing.

The finance director argued: “I very much doubt whether selling the Punch model is viable but I am not convinced that activity-based costing would tell us any more than the use of labour hours in assessing the viability of each product”.

The marketing director argued: “I am in the process of negotiating a major new contract with a motorcycle rental company for Discover model. For such a big order they will not pay our normal prices but we need to at least cover our incremental costs. I am not convinced that activity-based costing would achieve this as it merely averages costs for our entire production”.

The managing director argued: “I believe that activity-based costing would be an improvement but it still has its problems. For instance, if we carry out an activity many times surely we get better at it, and costs fall rather than remain constant. Similarly, some costs are fixed and do not vary either with labour hours or any other cost driver”.

The chairman argued: “I cannot see the problem. The overall profit for the company is the same no matter which method of allocating overheads we use. It seems to make no difference to me”.

REQUIRED:

  1. Calculate the total profit on each of Oriental Motors three types of product using each of the following methods to attribute overheads:
  1. The existing methods based upon labour hours.           
  2. Activity-based costing.                                                                
Answer #1

I. Overhead rate = Total overhead costs/Total direct labor hours

Total overhead costs = $2400000 + 6000000 + 3600000 = $12000000

Total direct labor hours = 200000 + 220000 + 80000 = 500000

Overhead rate = $12000000/500000 = $24 per direct labor hour

Discover Pulsar Punch
Sales 8000000 9600000 3200000
Less: Costs
Raw material cost 800000 960000 360000
Direct labor cost 1000000 1100000 400000
Overheads cost 4800000 5280000 1920000
Total cost 6600000 7340000 2680000
Total profit 1400000 2260000 520000

Working:

Discover Pulsar Punch
Sales =2000*4000 =1600*6000 =400*8000
Less: Costs
Raw material cost =2000*400 =1600*600 =400*900
Direct labor cost =200000*5 =220000*5 =80000*5
Overheads cost =200000*24 =220000*24 =80000*24

II.

Discover Pulsar Punch
Activity Activity Cost $ Cost Driver Volumes Activity Rate $ Cost Driver Usage Allocated Cost Cost Driver Usage Allocated Cost Cost Driver Usage Allocated Cost
Deliveries to retailers 2400000 250 9600 100 960000 80 768000 70 672000
Set-up costs 6000000 100 60000 35 2100000 40 2400000 25 1500000
Purchase orders 3600000 800 4500 400 1800000 300 1350000 100 450000
Total 12000000 4860000 4518000 2622000
Discover Pulsar Punch
Sales 8000000 9600000 3200000
Less: Costs
Raw material cost 800000 960000 360000
Direct labor cost 1000000 1100000 400000
Overheads cost 4860000 4518000 2622000
Total cost 6660000 6578000 3382000
Total profit 1340000 3022000 -182000